You have 0.5 ml of a patient's serum and you need to make a 1:100 dilutions of it in saline. How much saline would you add to your sample to accomplish that?

Please show work on how to solve. Thank you.
Answer #1

1:100 dilution means that 1 part of original (stock solution) is mixed with 99 parts of dilutent (or solvent) to make final volume of 100 part. 1:100 dilution can also be written as 1/100

Volume of original solution/total final volume = 1/100 (because given dilution ratio is 1/100)

  • Given volume of patient blood (original solution) = 0.5 ml

Therefore Total final volume required = 100 × 0.5 ml = 50 ml

  • Therefore amount of saline required = 50 - 0.5 = 49.5 ml

Feel free to comment if you have any problem in understanding this answer